OpenStudy (anonymous):

Need help ill give a medal Simplify the expression. 57 – 12 / 4 * 3

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

57-12 /4*3 = 45/12 = 15/4

OpenStudy (anonymous):

thanks